2. Suppose that Marissa is x years of age.
a. Write a mathematical expression for: The age of Marissa’s mother is twice the sum of Marissa’s age and 3.
b. How old is Marissa’s mother if Marissa is x = 18 years of age

Answer by macston(5194)

M=Mother's age
M=2(x+3)
For x=18yrs
M=2(x+3)=2(18yrs+3)=2(21yrs)=42 yrs ANSWER: Marissa's mother is 42 years old.